A shot fired with velocity u at elevation theta hits a mark at a distance R in the same horizontal plane. If there is no air resistance, show that (u2/g)f(theta).

1
Expert's answer
2021-01-28T13:41:17-0500

Maximum velocity will be u sin"\theta".

Now, equate it to maximum potential energy.

"m*g*h=(1\/2)mu^2"

"h = (1/2)u^2sin^2\theta/g" , here Sin "\theta" part is function of "\theta" .